OpenStudy (anonymous):

Re-asked: √x-2 + √x+3 ?

OpenStudy (anonymous):

Add like terms

OpenStudy (anonymous):

So it just becomes 2x +1?

OpenStudy (anonymous):



OpenStudy (anonymous):

Why not over it all, the sqrt is supposed to be covering the whole thing?

OpenStudy (anonymous):

So, do you mean: \[ \sqrt{x-2}+\sqrt{x+3} \]?

OpenStudy (anonymous):

Yeah

mathslover (mathslover):

\[\sqrt{x-2}+\sqrt{x+3}\]?

OpenStudy (anonymous):

You cannot simplify this further. Not directly, at least.

mathslover (mathslover):

square both sides by letting rhs be n

OpenStudy (anonymous):

Thats it...... nothing can be done. thats the simplified form

mathslover (mathslover):

squaring?

OpenStudy (anonymous):

But there has to be a way, this a question for my test review?

mathslover (mathslover):

square it

OpenStudy (anonymous):

No, you could do difference of squares, but you'd have to divide again. There's literally no other simplifications.

mathslover (mathslover):

hmn right

OpenStudy (anonymous):

The answer in the answer bank was [2,infinity), I think the end result should be the domain...?

OpenStudy (anonymous):

Okay, that makes more sense, then.

OpenStudy (anonymous):

We have that to make sure that \(\sqrt{x}\) is real, \(x>0\). Which means that: \[ x-2>0\implies\\ x>2 \]And: \[ x+3>0\implies\\ x>-3 \]Since, we have one more strict condition, we take the former, and so, we get \[ x\in[2,\infty) \]

OpenStudy (anonymous):

Oh I get it now, it has to be > than 0. Thanks a lot, really helped.

OpenStudy (anonymous):

Sure thing.